Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

Use the Henderson-Hasselbalch equation to perform the following calculations.

The

K

a

of acetic acid is 1.8

10

–5

.

a. Buffer A: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 0.5 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.

b. Buffer B: Calculate the mass of solid sodium acetate required to mix with

100.0 mL of 1.0 M acetic acid to prepare a pH 4 buffer. Record the mass in

your data table.
Answer #1

(a): Buffer-A:

Moles of acetic acid, CH3COOH = C*V = 0.5 mol/L * 100. mL * (1L/1000 mL) = 0.05 mol

Since this is a buffer solution, we have to use henderson equation

pH = pKa + log([CH3COONa] / [CH3COOH])

Since volume for both CH3COONa and CH3COOH are same, we can replace concentration by mole.

pH = pKa + log(moles of CH3COONa / moles of CH3COOH)

=> 4.0 = - log(1.8*10-5) + log(moles of CH3COONa / 0.05)

=> 4.0 = 4.745 + log(moles of CH3COONa / 0.05)

=> log(moles of CH3COONa / 0.05) = 4.0 - 4.745 = - 0.745

=> moles of CH3COONa / 0.05 = 10-0.745 = 0.180

=> moles of CH3COONa = 0.05 * 0.180

=> moles of CH3COONa = 0.0090 mol

=> mass of solid sodium acetate, CH3COONa required = 0.0090 mol * (82.0343 g/mol) = 0.74 g (Answer)

(b): Buffer-B:

Moles of acetic acid, CH3COOH = C*V = 1.0 mol/L * 100. mL * (1L/1000 mL) = 0.1 mol

Since this is a buffer solution, we have to use henderson equation

pH = pKa + log([CH3COONa] / [CH3COOH])

Since volume for both CH3COONa and CH3COOH are same, we can replace concentration by mole.

pH = pKa + log(moles of CH3COONa / moles of CH3COOH)

=> 4.0 = - log(1.8*10-5) + log(moles of CH3COONa / 0.1)

=> 4.0 = 4.745 + log(moles of CH3COONa / 0.1)

=> log(moles of CH3COONa / 0.1) = 4.0 - 4.745 = - 0.745

=> moles of CH3COONa / 0.1 = 10-0.745 = 0.180

=> moles of CH3COONa = 0.1 * 0.180

=> moles of CH3COONa = 0.0180 mol

=> mass of solid sodium acetate, CH3COONa required = 0.0180 mol * (82.0343 g/mol) = 1.48 g (Answer)